# PriRoAgg: Achieving Robust Model Aggregation with Minimum Privacy Leakage for Federated Learning

### v1

@@ IV Building Blocks of PriRoAgg
IV-A Key Agreement (KA)
We use the standard Key agreement (KA) and Authenticated encryption (AE) for user communications’ privacy and integrity. KA involves three algorithms and AE involves two algorithms.
@@ VI-A Security analysis
PriRoAgg EIFFeL
User communication $\mathcal{O}(\frac{Nd}{K})$ $\mathcal{O}(NTd)$
@@
    According to (12), the norm of $\tilde{\mathbf{x}}_{i}$ is constrained by $||\tilde{\mathbf{x}}_{i}||_{2}=1/\tilde{\omega}_{i}$. $\mathcal{S}$ samples $\tilde{\mathbf{x}}_{i}$ with a fixed $l_{2}$-norm, such that
@@ VII Experiments
We conduct experiments to demonstrate robust performance and computational overhead of PriRoAgg instantiations. We consider several targeted and untargeted attacks on two datasets. All experiments are simulated on a single machine using Intel(R) Xeon(R) Gold 5118 CPU @ 2.30GHz and four NVIDIA GeForce RTX 4090 GPUs.
@@ VII-A Setup
The untargeted and targeted attacks. For different types of attacks, we choose two attack strategies for each of trivial update poisoning, omniscient update poisoning, and backdoor attacks. For trivial update poisoning, the malicious update is manipulated by adding Gaussian noise [73] or scaling by constants [84]. For omniscient update poisoning, Min-Max and Min-Sum attack [46] are considered, where the attacker knows all benign updates to design its malformed update. For backdoor attacks, we consider image trojan attacks [85] with two different pattern types. We use the validation accuracy to demonstrate comparisons against other privacy-preserving protocols, and present the protocol’s overhead with different parameter settings.
@@
Figure 2: Performance of different attacks and defenses on Fashion-MNIST and CIFAR10. (a)-(f) are model poisoning attacks and (g)-(l) are backdoor attacks. Specifically, additive noise (a)(b), scaling (c)(d), Min-Max (e) and Min-Sum (f) attacks are demonstrated. (g)(h), (i)(j) and (k)(l) show PriRoAgg with RLR compared with EIFFeL, RoFL and EPPRFL, respectively. Sub-figures with the same attack share the same legend.
Baselines. We consider protocols with both robustness and privacy guarantees. BREA [32] and ByzSecAgg [33] ensure robustness using the Byzantine-resilient algorithm, multi-Krum [14], but both reveal pairwise distances to the server, thus violating aggregated privacy notion in our context. RoFL [31] and EIFFeL [30] provide theoretical guarantees for privacy in the same FL settings. They utilize the validation functions to ensure the updates’ robustness, with $L_{2}$ and $L_{\infty}$ norm bounding used as the baselines. We additionally use cosine similarity for EIFFeL’s implementation to present the effect of different choices of validation. Note that RoFL and EIFFeL’s validation functions rely on a public dataset and a given threshold; nevertheless, we compare PriRoAgg with them without any additional advantage. We also consider a two-server privacy-preserving protocol, EPPRFL, for backdoor attacks [36]. It uses the F&C algorithm adopted from [86] to filter malicious updates based on historical updating information.
Datasets. We use Fashion-MNIST [87], CIFAR-10 [88] to evaluate performance and test the end-to-end execution time. We assume the data is independent and identically distributed (i.i.d.) unless otherwise specified.
We choose $N=10$ for all Fashion-MNIST experiments and $N=40$ for all CIFAR-10 experiments, with $10\%-30\%$ malicious users. The security threshold is set to $T=\left\lceil 0.3N\right\rceil$ and $K=\left\lfloor 0.7N\right\rfloor-1$ to maximize efficiency while ensuring compliance with privacy guarantee. Each experiment is run for $10$ times and the average results are reported.
@@ VII-B Results
Performance. In Fig. 2 (a)-(f), the performance of the global model, which is measured by accuracy on the validation dataset, after applying different model poisoning attacks, is shown. PriRoAgg is consistently robust against model poisoning attacks. The multi-Krum algorithm demonstrates convergence in each scenario on Fashion-MNIST, but its accuracy is not as high as the other convergent baselines. For EIFFeL, its performance depends on the chosen validation predicate. For instance, EIFFeL with cosine similarity defends poorly against the Min-Max attack shown in (e) but succeeds in the Min-Sum attack in (f). RoFL performs well on Fashion-MNIST; however, it is no longer robust when CIFAR-10 is used. As CIFAR-10 serves as a more complex dataset, these baseline protocols become less effective. The multi-Krum algorithm diverges eventually, while RoFL experiences a drastic decrease around the $400$th iteration. EIFFeL exhibits relatively better performance with a different choice of validation, but it falls short in terms of accuracy and stability when compared to PriRoAgg.
@@
For backdoor attacks in (g)-(l) on Fig. 2, we evaluate the main task on the validation dataset and test the backdoor task on the poisoned validation dataset. As shown in (i) and (k), although RoFL and F&C manage to slow down the implantation process on Fashion-MNIST, both ultimately succumb to the backdoor. Compared to them, PriRoAgg with RLR performs significantly well against backdoor implantation on both datasets. It greatly supports the discussion in Section II-E about validation on the sole update not sufficiently supporting sophisticated model aggregation. The EIFFeL with cosine similarity has nearly no defensive capability against backdoor attacks, while RoFL with norm bounding performs slightly better. We consider multiple parameter settings for EPPRFL with F&C and find that F&C highly relies on its early-stage updates’ history; that is, if it cannot detect the malicious users in the first several rounds, it fails in the backdoor attack. Hence, F&C exhibits notable instability. As PriRoAgg allows the overall statistics of the updates, PriRoAgg with RLR robustly defends against the backdoor in both datasets, as well as achieving the best performance in the main task.
@@
Complexity results. We present the end-to-end runtime of PriRoAgg with the instantiations of RLR and RFA. The instantiation with RFA is presented with the default setting of $N=10$ and $d=1000$. To provide a comprehensive and fair comparison, we implement EIFFeL and use identical encoding and proving algorithms across protocols.
@@
Figure 3: The user and the server overhead breakdown of PriRoAgg per iteration. $N,d,T$ are the number of users, dimension of update, and maximum number of malicious users, respectively.
@@
Figure 4: The end-to-end overhead of PriRoAgg compared with EIFFeL. $N,d,T$ are the number of users, dimension of update, and maximum tolerance for malicious users, respectively.
@@
As shown in Fig. 3, the time against the number of users and the dimension of update follows the complexity analysis in Table  I. The time cost increases almost linearly as it is primarily dominated by variables $d$ and $N$. For time against $T$ with fixed $N$, the LCC parameter $K$ is chosen as the maximum, $N-T-1$, inducing a quadratic relation between time and $T$. In Fig.4, compared with EIFFeL, PriRoAgg benefits from higher encoding efficiency of LCC and repeated pattern circuit in SNIP. Consequently, PriRoAgg is superior in end-to-end efficiency, especially when $N$ is large since it allows for a larger $K$ for LCC. This advantage is particularly evident in Fig.4.

### v2

@@ IV Building Blocks of PriRoAgg
IV-A Key Agreement (KA) and Authenticated Encryption (AE)
We use the standard Key agreement (KA) and Authenticated encryption (AE) for user communication’s privacy and integrity. KA and AE involve three and two algorithms respectively.
@@ VI-A Security analysis
    According to (12), the norm of $\tilde{\mathbf{x}}_{i}$ is constrained by $||\tilde{\mathbf{x}}_{i}||_{2}=1/\tilde{\omega}_{i}$. $\mathcal{S}$ samples $\tilde{\mathbf{x}}_{i}$ with fixed $l_{2}$-norm, such that
@@
PriRoAgg EIFFeL
User communication $\mathcal{O}(\frac{Nd}{K})$ $\mathcal{O}(NTd)$
@@ VII Experiments
We conduct experiments to demonstrate robust performance and computational overhead of PriRoAgg instantiations. We consider several targeted and untargeted attacks on two datasets. All experiments are simulated on a single machine using Intel(R) Xeon(R) Gold 5118 CPU @ 2.30GHz and one NVIDIA GeForce RTX 4090 GPU. Our code is publicly available at this repository github.com/tardisblue9/PriRoAgg.
@@ VII-A Setup
The untargeted and targeted attacks. For different types of attacks, we choose various attack strategies among model update poisoning and backdoor attacks. Two trivial poisoning strategies are considered as baselines. The malicious users manipulate updates by adding Gaussian noise [80] or scaling by constants to their updates[90]. Moreover, omniscient update poisoning strategies are also used, such as Min-Max and Min-Sum attack [52], where the attacker knows all benign updates to design its malformed update. Two advanced poisoning strategies are also evaluated: PoisonFL [91] and Sine [92]. For backdoor attacks, we consider conventional image trojan attacks [93] with two different pattern types and advanced backdoor attack, IBA[94]. For performance comparisons, we use the global testing accuracy to evaluate the poisoning effectiveness, and then split the global test dataset into a clean test dataset and a backdoor test dataset to evaluate main task (main accuracy, MA $\uparrow$) and backdoor task accordingly (attack succuss rate, ASR $\downarrow$). For model poisoning, the main accuracy is also used for evaluation; a higher MA indicates better defense performance. Additionally, we present the protocol’s overhead with different parameter settings.
@@
Figure 2: Learning dynamics of different attacks and defenses on Fashion-MNIST and CIFAR10. (a)-(f) are model poisoning attacks and (g)-(l) are backdoor attacks. Specifically, additive noise (a)(b), scaling (c)(d), Min-Max (e) and Min-Sum (f) attacks are demonstrated. (g)(h), (i)(j) and (k)(l) show PriRoAgg with RLR compared with EIFFeL, RoFL and EPPRFL, respectively. Sub-figures with the same attack share the same legend.
Baselines. For defenses, we consider protocols that attempts to achieve robustness and privacy guarantees. BREA [32] and ByzSecAgg [33] ensure robustness using the Byzantine-resilient algorithm, multi-Krum [14]. RoFL [31] and EIFFeL [30] utilize the validation functions to ensure the updates’ robustness, where $L_{2}$ and $L_{\infty}$ norm bounding are used as the baselines. We additionally use historical cosine similarity for EIFFeL’s implementation to present the effect of different choices of validation functions. Note that RoFL and EIFFeL’s validation functions rely on a public dataset and a given threshold, nevertheless, we allow their advantages with the dataset, which is isolated from the training dataset. For similarity and rank schemes in Trust-ZKFL [47], they use Multi-Krum and SignGuard [71], respectively. Consequently, similarity-based Trust-ZKFL corresponds to the Multi-Krum case mentioned before, while rank-based Trust-ZKFL is presented for comparison. For scarcity of backdoor-resilient privacy-preserving protocols in one-server setting, we consider another two-server protocol, EPPRFL, for backdoor attacks [36]. It uses the F&C algorithm adopted from [95] to filter malicious updates based on historical updating information.
Datasets. We use Fashion-MNIST (FMNIST) [96], CIFAR-10 [97] to evaluate performance and test the end-to-end execution time. We assume the data is independent and identically distributed (i.i.d.) unless otherwise specified.
We choose $N=10$ for all Fashion-MNIST experiments and $N=40$ for all CIFAR-10 experiments, with $10\%-30\%$ malicious users. The security threshold is set to $T=\left\lceil 0.3N\right\rceil$ and $K=\left\lfloor 0.7N\right\rfloor-1$ to maximize efficiency while ensuring compliance with privacy guarantee. Each experiment is run for $10$ times and the average results are reported. The number of global epoch is set to $500$.
@@ VII-B Results
TABLE II: Model Poisoning Performance
Fashion-MNIST CIFAR-10
No Att. Gaussian Scale MinMax MinSum PoisonFL Sine No Att. Gaussian Scale MinMax MinSum PoisonFL Sine
No Defence 93.8 10.0 10.0 10.0 10.0 10.0 10.0 79.5 10.0 10.0 10.0 10.0 10.0 10.0
Multi-Krum - 89.6 89.5 89.5 89.9 89.8 89.4 - 13.3 14.7 14.1 14.0 12.6 15.4
EIFFel ($L_{2}$) - 92.8 93.3 68.5 92.9 93.1 92.5 - 77.6 77.9 75.5 75.6 77.2 77.9
EIFFel (cosine) - 10.0 10.0 68.9 92.6 10.0 10.0 - 10.0 10.0 10.0 10.0 10.0 10.0
RoFL ($L_{\infty}$) - 93.3 93.4 92.3 92.5 93.2 92.6 - 11.8 27.8 18.8 19.0 17.4 18.5
SignGuard - 93.5 91.1 71.0 80.5 93.3 10.0 - 79.5 78.8 10.0 11.3 78.7 10.0
PriRoAgg (ours) - 93.6 93.6 92.6 93.1 92.2 92.9 - 79.5 79.5 79.2 79.3 78.6 79.1
@@
Performance. We present the final accuracy for every attack-defense pair in Table  II,  III. We also notice the effectiveness of defenses changes throughout the training process, hence demonstrate the learning dynamics of selected experiments in Fig. 2. The performance is measured by the testing accuracy of the global model and is presented as a percentage. For model poisoning attacks, PriRoAgg consistently yields robust results close to the no-attack baseline. The multi-Krum algorithm demonstrates convergence on FMNIST, although its accuracy is not as high as that of other convergent baselines. The performance of EIFFeL depends on the chosen validation predicate. On FMNIST, EIFFeL with cosine similarity poorly defends against all attacks except the Min-Sum attack, and EIFFeL with $L_{2}$-norm fails in the Min-Max attack. RoFL generally performs well on FMNIST; however, it is no longer robust when using CIFAR-10. Trust-ZKFL with SignGuard experiences a performance degradation in omniscient attacks such as Min-Max, Min-Sum, and Sine attacks. As CIFAR-10 is a more complex dataset, multi-Krum, EIFFeL with cosine similarity, and RoFL fail to prevent model poisoning. Instead of failing at the beginning of the training stage, RoFL experiences a drastic decrease around the $400$th iteration in Fig. 2(b). This behavior is pervasive when the learning dynamics change the distribution of updates. EIFFeL exhibits relatively good performance with a correct choice of validation predicate, but it falls short in terms of accuracy and stability compared to PriRoAgg. Trust-ZKFL with SignGuard has comparable performance in simple attacks and PoisonFL attack, but PriRoAgg consistently perform well in every attack.
TABLE III: Backdoor Performance
FMNIST CIFAR10
Trojan A Trojan B IBA
MA ()($\uparrow$)( ↑ ) ASR ()($\downarrow$)( ↓ ) MA ASR MA ASR
No Defence 93.2 94.4 79.5 62.0 79.4 61.9
EIFFel ($L_{2}$) 93.1 100.0 79.0 38.9 78.0 27.5
EIFFel (cosine) 93.4 100.0 50.1 40.7 22.1 23.0
RoFL ($L_{\infty}$) 92.6 100.0 74.9 42.7 59.2 25.4
F&C 93.3 100.0 50.5 46.2 37.4 24.3
PriRoAgg (ours) 92.2 0.0 78.7 13.1 78.9 12.4
@@
For backdoor attacks, we evaluate the main task on the clean test dataset and the backdoor task on the poisoned test dataset. The learning dynamics are presented in (g)-(l) on Fig. 2 and the final results are reported in Table III. Generally, the compared defenses are effective in the early stages but fail in the long term. As shown in (i) and (k), although RoFL and F&C manage to slow down the implantation process on Fashion-MNIST, both ultimately succumb to the backdoor. Compared to them, PriRoAgg with RLR performs significantly well against backdoor implantation on both datasets. It greatly supports the discussion in Section II-E about validation on the sole update not sufficiently supporting sophisticated model aggregation. The EIFFeL with cosine similarity has nearly no defensive capability against backdoor attacks, while EIFFeL and RoFL with norm bounding performs slightly better. However, they significantly reduces the main task performance when facing advanced attacks such as IBA. We consider multiple parameter settings for EPPRFL with F&C and notice that F&C highly relies on its early-stage updates’ history; that is, if it cannot detect the malicious users in the first several rounds, it fails in the backdoor attack. Hence, F&C exhibits notable instability shown in (l). In comparison, by allowing the overall statistics to aid aggregation, PriRoAgg robustly defends against the backdoor in both backdoor attacks, while achieving the comparable performance in the main task as baseline.
@@
Complexity results. We present the end-to-end runtime of PriRoAgg with the instantiations of RLR and RFA. The instantiation with RFA is presented with the default setting of $N=10$ and $d=1000$. To provide a comprehensive and fair comparison, we implement EIFFeL and use identical encoding, decoding and ZKP algorithms across protocols.
@@
Figure 3: The user and the server overhead breakdown of PriRoAgg per iteration. $N,d,K$ are the number of users, dimension of update, and LCC parameter, respectively.
@@
Figure 4: The end-to-end overhead of PriRoAgg compared with EIFFeL. $N,d,K$ are the number of users, dimension of update, and LCC parameter, respectively.
@@
As shown in Fig. 3, the time against the number of users and the dimension of update follows the complexity analysis in Table  I. The time cost increases linearly as it is primarily dominated by variables $d$ and $N$.For different LCC encoding parameter $K$ with fixed $T,N$ satisfying $K<N-T$, the larger LCC parameter $K$ shows higher encoding efficiency, where EIFFeL is approximately $K=1$ case. In Fig.4, compared with EIFFeL, PriRoAgg benefits from higher encoding efficiency of LCC and repeated pattern circuit in improved SNIP. In summary, PriRoAgg is superior in end-to-end efficiency, especially when $N$ is large allowing larger $K$ for LCC. This advantage is particularly evident in Fig.4.
@@ VIII Conclusion
Acknowledgments
This work is supported in part by the Fundamental Research Funds for the Central Universities (Grant No. 2242025K30025). The work of T. Jahani-Nezhad and G. Caire was supported by the Gottfried Wilhelm Leibniz Prize 2021, awarded by the German Research Foundation (DFG).